Three bars are connected with frictionless pins to form the frame shown. The frame is loaded with a force of P = 670 lb. TheCA kab c dRp, = RD, = (C)
Answer #1

Answer Draw free body diagrams FBD 0 ABC FBD oH BDF FBD 4 CDE - Rad Rex RO 67016 Ky RAY Here, o = tani (+) = 63.43 (a) RB=? fAAA { Fy=0 Rg Cofo + Roy -670 =0 @ → Roy = 670 + R3 CO3 (63.43) = 670 +1395.2685(63.43) > Roy = 1792:55W Aus •Rex, Rey Take→ Rey - Roy - Rey = 1792-55-476-2 3 =1316.35 Rey = 1316-35 16 ans EME=0 → Rey(12.1) - Rcq(12.1 tano) - Roy(+) –RDX (7tang) =0→ Rex = Rcx + RDX = 455+ (-12A7-87) = -79266 . Rex = -792 lb Ay Rey = 476.26b Aus from eq @ (f) RA, RC, RO, RE = 81619 Db RA